OpenStudy (anonymous):

find a general solution for the system of differential equations: dx/dt=-7x+10z; dy/dt=9x+2y-11z; dz/dt=-8x+11z

OpenStudy (anonymous):

use it as implicit diffrentation...

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

i think it was, dz/dx=dz/dt.dt/dx

OpenStudy (anonymous):

x'=-7x+10z y'=9x+2y-11z z'=-8x+11z

OpenStudy (anonymous):

(D+7)x+ 0y -10z =0 -9x +(D-2)y +11z =0 8x + 0 y +(D-11)z =0